OpenStudy (anonymous):

A particle is moving along the curve y=2*sqrt(5x+11). As the particle passes through the point (5,12) , its x -coordinate increases at a rate of 2 units per second. Find the rate of change of the distance from the particle to the origin at this instant. I'm having trouble getting dy/dt and y

OpenStudy (anonymous):

distance = sqrt (x^2 +y^2) =sqrt (x^2 +(2*sqrt(5x+11))^2) differentiate with respect to t...

OpenStudy (anonymous):

Sounds like a related rates situation, right?

OpenStudy (anonymous):

plug in dx/dt and x...

OpenStudy (anonymous):

yes related rates

OpenStudy (anonymous):

...simplify a bit and you're done:)

OpenStudy (anonymous):

OK, good, I thought it was going to get all parametric and crap. Looks like @Algebraic! laid the course for you well.

OpenStudy (anonymous):

why would the y coordinate not be 12?

OpenStudy (anonymous):

you mean just plug in y=12 into sqrt (x^2 +y^2) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that would be easy but wrong:)

OpenStudy (anonymous):

ok I calculated that y=72, so I'll try this......

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

when finding y, do I plug in 5 for x for 2sqrt(5x+11)?

OpenStudy (anonymous):

into 2sqrt(5x+11), i mean....

OpenStudy (anonymous):

that's already been done for you. y=12

OpenStudy (anonymous):

you should really try finding the derivative of: sqrt (x^2 +(2*sqrt(5x+11))^2) that's how you complete this problem.

OpenStudy (anonymous):

simplify it a little and differentiate. use the chain rule. the derivative of 'x' leave as dx/dt

OpenStudy (anonymous):

alright i'm going to try and start over

OpenStudy (anonymous):

when i try to differentiate y=2sqrt(5x+11) i get 5(5x+11)^(-1/2) using the chain rule. This has to be what I keep getting wrong.

OpenStudy (noelgreco):

You need to differentiate that function WRT time because you need to know dy/dt

OpenStudy (anonymous):

\[\frac{ d }{ dt } \sqrt{ (x^2 +(2*\sqrt(5x+11))^2)}) = \frac{ d }{ dt } \sqrt{ (x^2 +20x+44})\] \[ \frac{ d }{ dt } \sqrt{ (x^2 +20x+44}) =\frac{ (2x+20)\frac{ dx }{ dt } }{2 \sqrt{ (x^2 +20x+44}) }\] plug in values of x and dx/dt

OpenStudy (anonymous):

the computer system must be wrong, i got that answer too. thank you for your patience and help

OpenStudy (anonymous):

what did you get for the final answer?

OpenStudy (anonymous):

2.307692308

OpenStudy (anonymous):

30/13

OpenStudy (anonymous):

it can be entered as a fraction or in radians, so yes, samething

OpenStudy (noelgreco):

\[y=2(5x+11)^{\frac{ 1 }{ 2 }}\] \[\frac{ dy }{ dt }=5(5x+11)^{\frac{ -1 }{ 2 }}\frac{ dx }{ dt }\] dx/dt = 2 (given), x=5 Solve for dy/dt. Then:  \[x ^{2}+y ^{2}=g ^{2}\]

OpenStudy (noelgreco):

Differentiate WRT time, solve for dg/dt, plug in values, and have a cigar.

OpenStudy (anonymous):

Thank you!